# Generating Molecular Conformer Fields

### v2

@@ 4 Method
MCF is a diffusion generative model that captures distributions over conformer fields. We are given observations in the form of an empirical distribution $f_{0}\sim q(f_{0})$ over fields where a field $f_{0}:\mathcal{G}\rightarrow\mathbb{R}^{3}$ maps vertices $v\in\mathcal{G}$ on a molecular graph $\mathcal{G}$ to 3D space $\mathbb{R}^{3}$. As a result, latent variables $f_{1:T}$ are also functions on graphs that can be continuously evaluated.
where $\mathbf{\epsilon}_{c}\sim\mathcal{N}({\bm{0}},\mathbf{I})$ is a noise vector of the appropriate size. We now turn to the task of formulating a score network for fields. The score network needs to take as input the context set (i.e.   the field parametrization), and needs to accept being evaluated continuously in $\mathcal{G}$. We do this by sampling a query set of vertex-signal pairs $\{\varphi(v_{q}),{\bm{y}}_{(q,0)}\}$. Equivalently to the context set, we row-wise stack query pairs and denote the resulting matrix as ${\mathbf{Q}}_{0}~{}=~{}[\varphi(\mathcal{V}_{q}),~{}{\mathbf{Y}}_{(q,0)}]$. Note that the forward diffusion process is equivalently defined for both context and query sets:
Using the explicit field characterization and the score field network, we obtain the training and inference procedures in Alg. 1 and Alg. 2, respectively, which are accompanied by illustrative examples of sampling a field encoding a Gaussian mixture model over the manifold (i.e.   the bunny). For training, we uniformly sample context and query sets from $f_{0}\sim\mathrm{Uniform}(q(f_{0}))$ and only corrupt their signal using the forward process in Eq. equation 1 and Eq. equation 2. We train the score field network $\epsilon_{\theta}$ to denoise the signal portion of the query set, given the context set. During sampling, to generate a conformer fields $f_{0}\sim p_{\theta}(f_{0})$ we first define a query set ${\mathbf{Q}}_{T}~{}=~{}[\varphi(\mathcal{V}_{q}),~{}{\mathbf{Y}}_{(q,T)}\sim~{}\mathcal{N}({\bm{0}},~{}{\mathbf{I}})]$ of random values to be de-noised. We set the context set to be a random subset of the query set. We use the context set to denoise the query set and follow ancestral sampling as in the vanilla DDPM (Ho et al., 2020). Note that during inference the eigen-function representation $\varphi(v)$ of the context and query sets does not change, only their corresponding signal value (e.g.  their 3D position).
@@ 4.1 Score Field Network $\epsilon_{\theta}$
In MCF, the score field’s design space covers all architectures that can process irregularly sampled data, such as Transformers (Vaswani et al., 2017) and their corresponding Graph counterparts (Maskey et al., 2022; Sharp et al., 2022; He et al., 2022; Dwivedi et al., 2020) which have recently gained popularity in the supervised learning setting. The score field network $\epsilon_{\theta}$ is primarily implemented using PerceiverIO (Jaegle et al., 2022), an effective transformer encoder-decoder architecture. The PerceiverIO was chosen due to its efficiency in managing large numbers of elements in the context and query sets, as well as its natural ability to encode interactions between these sets using attention. Figure 4 demonstrates how these sets are used within the PerceiverIO architecture. To elaborate, the encoder maps the context set into latent arrays (i.e.  a group of learnable vectors) through a cross-attention layer, while the decoder does the same for query set. For a more detailed analysis of the PerceiverIO architecture refer to (Jaegle et al., 2022).The time-step $t$ is incorporated into the score computation by concatenating a positional embedding representation of $t$ to the context and query sets.
Figure 4: Interaction between context and query pairs in the PerceiverIO architecture. Context pairs $\mathbf{C}_{t}$ attend to a latent array of learnable parameters via cross attention. The latent array then goes through several self attention blocks. Finally, the query pairs $\mathbf{Q}_{t}$ cross-attend to the latent array to produce the final noise prediction $\hat{\epsilon_{q}}$.
@@ 5 Experiments
We use two popular datasets: GEOM-QM9 (Ruddigkeit et al., 2012) and GEOM-DRUGS (Ruddigkeit et al., 2012). Datasets are preprocessed as described in GeoMol (Ganea et al., 2021). We split them randomly based on molecules into train/validation/test (80%/10%/10%). At the end, for each dataset, we sample 1000 random test molecules as the final test set. Thus, the splits contain 106586/13323/1000 and 243473/30433/1000 molecules for GEOM-QM9 and GEOM-DRUGS, resp. We follow the exact same training splits for all baselines. In the Appendix A.3 we provide additional experiments that validate the design choices for the score network architecture, as well as empirically validating the chemical properties of generated conformers.
@@ 5.1 GEOM-QM9
Following the standard setting for molecule conformer prediction we use the GEOM-QM9 dataset (Ruddigkeit et al., 2012) which contains $\sim 130$K molecules ranging from 3 to 29 atoms. We report our results in Tab. 1 and compare with CGCF (Xu et al., 2021a), GeoDiff (Xu et al., 2022), GeoMol (Ganea et al., 2021) and Torsional Diffusion (Jing et al., 2022). Note that both GeoMol (Ganea et al., 2021) and Torsional Diffusion (Jing et al., 2022) make strong assumptions about the geometric structure of molecules and model domain-specific characteristics (i.e.  torsional angles of rotatable bonds). In contrast, MCF simply models the distribution of 3D coordinates of atoms without making any assumptions about the underlying structure. Finally we report the same metrics as Torsional Diff. (Jing et al., 2022) to compare the generated and ground truth conformer ensembles: Average Minimum RMSD (AMR) and Coverage. These metrics are reported both for precision, measuring the accuracy of the generated conformers, and recall, which measures how well the generated ensemble covers the ground-truth ensemble (details about metrics can be found in Appendix A.2.4). We generate $2K$ conformers for a molecule with $K$ ground truth conformers.
We report results in Tab. 1, showing that MCF outperforms previous approaches by a substantial margin. In addition, it is important to note that MCF is a general approach for learning functions on graphs that does not make any assumptions about the intrinsic geometric factors important in conformers like torsional angles. This makes MCF simpler to implement and applicable to other settings in which intrinsic geometric factors are not known or expensive to compute.
@@
Coverage $\uparrow$ AMR $\downarrow$ Coverage $\uparrow$ AMR $\downarrow$
@@
CGCF 69.47 96.15 0.425 0.374 38.20 33.33 0.711 0.695
GeoDiff 76.50 100.00 0.297 0.229 50.00 33.50 0.524 0.510
GeoMol 91.50 100.00 0.225 0.193 87.60 100.00 0.270 0.241
Torsional Diff. 92.80 100.00 0.178 0.147 92.70 100.00 0.221 0.195
MCF (ours) 96.57 100.00 0.107 0.072 94.60 100.00 0.130 0.084
Table 1: Molecule conformer generation results on GEOM-QM9. MCF obtains better results than the state-of-the-art Torsional Diffusion (Jing et al., 2022), without making any explicit assumptions about the geometric structure of molecules (i.e.  without modeling torsional angles).
@@ 5.2 GEOM-DRUGS
To test the capacity of MCF to deal with larger molecules we also report experiments on GEOM-DRUGS, the largest and most pharmaceutically relevant part of the GEOM dataset (Axelrod & Gomez-Bombarelli, 2022) — consisting of 304k drug-like molecules (average 44 atoms). We report our results in Tab. 2 and compare with GeoDiff (Xu et al., 2022), GeoMol (Ganea et al., 2021) and Torsional Diffusion (Jing et al., 2022). Note again that both GeoMol (Ganea et al., 2021) and Torsional Diffusion (Jing et al., 2022) make strong assumptions about the geometric structure of molecules and model domain-specific characteristics like torsional angles of bonds. In contrast, MCF simply models the distribution of 3D coordinates of atoms without making any assumptions about the underlying structure.
Results on Tab. 2 are where we see how MCF outperforms strong baseline approaches by substantial margins. MCF achieves better performance than previous state-of-the-art model Torsional Diffusion by approximately 10% on both recall and precision. This indicates that our proposed MCF not only generates high-quality conformers that are close with ground truth but also covers a wide variety of conformers in the ground truth. In addition, it is important to note that MCF does not make any assumptions about the intrinsic geometric factors important in conformers like torsional angles and thus is simpler to train (e.g. it does not require a local substructure prediction model).
@@
Coverage $\uparrow$ AMR $\downarrow$ Coverage $\uparrow$ AMR $\downarrow$
@@
GeoDiff 42.10 37.80 0.835 0.809 24.90 14.50 1.136 1.090
GeoMol 44.60 41.40 0.875 0.834 43.00 36.40 0.928 0.841
Torsional Diff. 72.70 80.00 0.582 0.565 55.20 56.90 0.778 0.729
MCF (ours) 81.62 89.22 0.468 0.438 61.63 62.50 0.705 0.650
Table 2: Molecule conformer generation results on GEOM-DRUGS. MCF obtains comparable or better results than the state-of-the-art Torsional Diffusion (Jing et al., 2022), without making any explicit assumptions about the geometric structure of molecules (i.e.  without modeling torsional angles).
In addition, in Fig. 5 we show a breakdown of the performance on GEOM-DRUGS of MCF vs. Torsional diffusion (Jing et al., 2022) as a function of the threshold distance, as well as a function of the number of atoms in molecules. MCF outperforms Torsional Diffusion across the full spectrum of threshols in both recall and precision. When looking at the break-down AMR on different number of atoms in Fig. 5(c), MCF also demonstrates its superior performance in most cases. It is indicated that MCF better captures the fine intrinsic geometric structure of conformers.
@@
Figure 5: (a) Recall coverage metric as a function of the threshold distance. MCF outperforms Torsional Diffusion across the full spectrum of thresholds. (b) Precision coverage as a function of the threshold distance. Still MCF outperforms Torsional Diffusion across the full spectrum of thresholds. (c) Minimum averaged RMSD (lower is better) of recall and precision as a function of the number of atoms in molecules.
5.2.1 Generalization to GEOM-XL
We now turn to the task of evaluating how well a model trained on GEOM-DRUGS transfers to unseen molecules with large numbers of atoms. As proposed in Torsional Diff. (Jing et al., 2022) we use the GEOM-XL dataset, which is a subset of GEOM-MoleculeNet that contains all species with more than 100 atoms, which is a total of 102 molecules. Note that this evaluation not only tests the capacity of models to generalize to molecules with large number of atoms but also serves as an out-of-distribution generalization experiment.
In Tab. 3 we report AMR for both precision and recall and compare with GeoDiff (Xu et al., 2022), GeoMol (Ganea et al., 2021) and Torsional Diffusion (Jing et al., 2022). In particular, when taking the numbers directly from (Jing et al., 2022), MCF achieves better or comparable performance than Torsional Diffusion. Further, in running the checkpoint provided by Torsional Diffusion and following their validation process we found that 25 molecules failed to be generated, this is due to the fact that Torsional Diffusion generates torsional angles conditioned on the molecular graph $\mathcal{G}$ and the local structures obtained from RDKit. In some cases, RDKit can fail to find local structures and Torsional Diffusion cannot generate conformers in this case. In our experiments with the same 77 molecules in GEOM-XL from our replica, MCF surpasses Torsional Diffusion by a large margin. The results highlight the generalizability of MCF to large and complex molecules, which may shed a light on pre-training molecular conformer generation model.
@@ 5.2.1 Generalization to GEOM-XL
MCF 1.77 1.59 2.62 2.37 77
Table 3: Generalization results on GEOM-XL. MCF obtains comparable results to Torsional Diffusion (Jing et al., 2022), without making any explicit assumptions about the geometric structure of molecules.
5.2.2 Ensemble property prediction
To fully assess the quality of generated conformers we also compute chemical property resemblance between the synthesized and the authentic ground truth ensembles. We select a random group of 100 molecules from the GEOM-DRUGS and produce a minimum of $2K$ and a maximum of $32$ conformers for each molecule following (Jing et al., 2022). Subsequently, we undertake a comparison of the Boltzmann-weighted attributes of the created and the true ensembles. To elaborate, we calculate the following characteristics using xTB (as documented by (Bannwarth et al., 2019)): energy ($E$), dipole moment ($\mu$), the gap between HOMO and LUMO ($\Delta\epsilon$), and the lowest possible energy, denoted as $E_{\text{min}}$. Since we don’t have the access to the exact subset of DRUGS used in (Jing et al., 2022), we randomly pick three subsets and report the averaged and standard deviation over three individual runs with different random seeds. The results are listed in Tab. 4. Our model achieves the lowest error on $E_{\text{min}}$ when compared with other baselines, which demonstrates that MCF is succeeds at generating stable conformers that are very close to the ground states. This could root from the fact that MCF doesn’t rely on rule-based cheminfomatics methods and the model learns to better model stable conformers from data. Besides, MCF achieves competitive performance on $\mu$ and $\Delta\epsilon$. However, the error of $E$ is high compared to the rest of approaches, meaning that though MCF performs well in generating samples close to ground states, it may also generate conformers with high energy that are not plausible in the dataset.
$E$ $\mu$ $\Delta\epsilon$ $E_{\text{min}}$
OMEGA 0.68 0.66 0.68 0.69
GeoDiff 0.31 0.35 0.89 0.39
GeoMol 0.42 0.34 0.59 0.40
Torsional Diff. 0.22 0.35 0.54 0.13
MCF 0.66$\pm$0.05 0.26$\pm$0.05 0.64$\pm$0.06 0.02$\pm$0.00

### v3

@@ 4 Method
MCF is a diffusion generative model that captures distributions over conformer fields. We are given observations in the form of an empirical distribution $f_{0}\sim q(f_{0})$ over fields where a field $f_{0}:\mathcal{G}\rightarrow\mathbb{R}^{3}$ maps vertices $v\in\mathcal{G}$ on a molecular graph $\mathcal{G}$ to 3D space $\mathbb{R}^{3}$.
@@
where $\mathbf{\epsilon}_{c}\sim\mathcal{N}({\bm{0}},\mathbf{I})$ is a noise vector of the appropriate size. We now turn to the task of formulating a score network for fields. The score network needs to take as input the context set (i.e.   the field parametrization), and needs to accept being evaluated for any point in $\mathcal{G}$. We do this by sampling a query set of vertex-signal pairs $\{\varphi(v_{q}),{\bm{y}}_{(q,0)}\}$. Equivalently to the context set, we row-wise stack query pairs and denote the resulting matrix as ${\mathbf{Q}}_{0}~{}=~{}[\varphi(\mathcal{V}_{q}),~{}{\mathbf{Y}}_{(q,0)}]$. Note that the forward diffusion process is equivalently defined for both context and query sets:
Using the explicit field characterization and the score field network, we obtain the training and inference procedures in Alg. 1 and Alg. 2, respectively, which are accompanied by illustrative examples of sampling a conformer field. For training, we uniformly sample context and query sets from $f_{0}\sim\mathrm{Uniform}(q(f_{0}))$ and only corrupt their signal using the forward process in Eq. equation 1 and Eq. equation 2. We train the score field network $\epsilon_{\theta}$ to denoise the signal portion of the query set, given the context set. During sampling, to generate a conformer fields $f_{0}\sim p_{\theta}(f_{0})$ we first define a query set ${\mathbf{Q}}_{T}~{}=~{}[\varphi(\mathcal{V}_{q}),~{}{\mathbf{Y}}_{(q,T)}\sim~{}\mathcal{N}({\bm{0}},~{}{\mathbf{I}})]$ of random atom positions to be de-noised. We set the context set to be a random subset of the query set. We use the context set to denoise the query set and follow ancestral sampling as in the vanilla DDPM (Ho et al., 2020). Note that during inference the eigen-function representation $\varphi(v)$ of the context and query sets does not change, only their corresponding signal value (e.g.  their 3D position).
@@ 4.1 Score Field Network $\epsilon_{\theta}$
In MCF, the score field’s design space covers all architectures that can process irregularly sampled data, such as Transformers (Vaswani et al., 2017) and their corresponding Graph counterparts (Maskey et al., 2022; Sharp et al., 2022; He et al., 2022; Dwivedi et al., 2020) which have recently gained popularity in the supervised learning setting. The score field network $\epsilon_{\theta}$ is primarily implemented using PerceiverIO (Jaegle et al., 2022), an effective Transformer encoder-decoder architecture. A PerceiverIO is chosen due to its nature of a general-purposed architecture that can handle data of a wide variety domains. It provides a suitable test bed for evaluating how well models without domain-specific inductive bias (e.g. equivariance) perform in solving scientific problems (e.g. molecular conformer generation as investigated in this work). PerceiverIO encodes interactions between elements in sets using attention, which has been demonstrated to be scalable in many previous works (Brown et al., 2020). Fig. 1 demonstrates how these sets are used within the PerceiverIO architecture. To elaborate, the encoder maps the context set into latent arrays (i.e.  a group of learnable vectors) through a cross-attention layer, while the decoder does the same for query set. For a more detailed analysis of the PerceiverIO architecture refer to (Jaegle et al., 2022).The time-step $t$ is incorporated into the score computation by concatenating a positional embedding representation of $t$ to both context and query sets.
@@ 5 Experiments
We use two popular datasets: GEOM-QM9 and GEOM-DRUGS (Axelrod & Gomez-Bombarelli, 2022). Datasets are preprocessed and split as described in Ganea et al. (2021). We deploy PerceiverIO with small (S), base (B) and large (L) sizes, which contain 13M, 64M and 242M parameters respectively. More implementation details can be found in Appendix A.2. We provide additional experiments that validate the design choices for the score network architecture, as well as empirically validating the chemical properties of generated conformers in the Appendix A.3.
@@ 5.1 GEOM-QM9
COV $\uparrow$ AMR $\downarrow$ COV $\uparrow$ AMR $\downarrow$
@@
CGCF 69.5 96.2 0.425 0.374 38.2 33.3 0.711 0.695
GeoDiff 76.5 100.0 0.297 0.229 50.0 33.5 0.524 0.510
GeoMol 91.5 100.0 0.225 0.193 87.6 100.0 0.270 0.241
Tor. Diff. 92.8 100.0 0.178 0.147 92.7 100.0 0.221 0.195
MCF 95.0 100.0 0.103 0.044 93.7 100.0 0.119 0.055
Table 1: Molecule conformer generation results on GEOM-QM9. MCF obtains better results than the state-of-the-art baselines.
Following the standard setting for molecule conformer prediction we use the GEOM-QM9 dataset which contains $\sim 130$K molecules ranging from 3 to 29 atoms. We report our results with base size model (i.e.  MCF-B) in Tab. 1 and compare with CGCF (Xu et al., 2021a), GeoDiff (Xu et al., 2022), GeoMol (Ganea et al., 2021) and Torsional Diff. (Jing et al., 2022). Note that all baselines make strong assumptions about the geometric structure of molecules. They either develop equivariant diffusion process (Xu et al., 2022) or model domain-specific characteristics like interatomic distances (Xu et al., 2021a) and torsional angles of rotatable bonds (Ganea et al., 2021; Jing et al., 2022). In contrast, MCF simply models the distribution of 3D coordinates of atoms without making any assumptions about the underlying structure. Finally we report the same metrics as Torsional Diff. (Jing et al., 2022) to compare the generated and ground truth conformer ensembles: average minimum RMSD (AMR) and coverage (COV). These metrics are reported both for precision, measuring the accuracy of the generated conformers, and recall, measuring how well the generated ensemble covers the ground-truth ensemble (details about metrics can be found in Appendix A.2.4). We generate $2K$ conformers for a molecule with $K$ ground truth conformers.
Tab. 1 shows that MCF outperforms previous approaches by a substantial margin. In addition, it is important to note that MCF is a general approach for learning functions on graphs that does not make any assumptions about the intrinsic geometric factors important in conformers like torsional angles. This makes MCF simpler to implement and applicable to other settings in which intrinsic geometric factors are not known or expensive to compute.
@@ 5.2 GEOM-DRUGS
COV $\uparrow$ AMR $\downarrow$ COV $\uparrow$ AMR $\downarrow$
@@
GeoDiff 42.1 37.8 0.835 0.809 24.9 14.5 1.136 1.090
GeoMol 44.6 41.4 0.875 0.834 43.0 36.4 0.928 0.841
Tor. Diff. 72.7 80.0 0.582 0.565 55.2 56.9 0.778 0.729
MCF-S 79.4 87.5 0.512 0.492 57.4 57.6 0.761 0.715
MCF-B 84.0 91.5 0.427 0.402 64.0 66.2 0.667 0.605
MCF-L 84.7 92.2 0.390 0.247 66.8 71.3 0.618 0.530
Table 2: Molecule conformer generation results on GEOM-DRUGS. MCF surpasses state-of-the-art baselines by large margin.
@@
Figure 4: (a) Recall coverage and (b) precision coverage as a function of the threshold distance. MCF outperforms Torsional Diff. across the full spectrum of thresholds. (c) Averaged AMR of recall and precision as a function of the number of atoms in molecules.
To test the capacity of MCF to deal with larger molecules we also report experiments on GEOM-DRUGS, the largest and most pharmaceutically relevant part of the GEOM dataset (Axelrod & Gomez-Bombarelli, 2022) — consisting of 304k drug-like molecules (average 44 atoms). We report our results in Tab. 2 and compare with GeoDiff (Xu et al., 2022), GeoMol (Ganea et al., 2021) and Torsional Diff. (Jing et al., 2022). Note again that all baseline approaches make strong assumptions about the geometric structure of molecules and model domain-specific characteristics like torsional angles of bonds. MCF simply models the distribution of 3D coordinates of atoms without making any assumptions about the underlying structure.
Results on Tab. 2 are where we see MCF outperforms strong baseline approaches by substantial margins. All MCF models achieve better performance than previous state-of-the-art Torsional Diff. model. On both recall and precision, MCF of small size (MCF-S) surpasses Torsional Diff. by approximately 15%. This indicates that our proposed MCF not only generates high-quality conformers that are close with ground truth but also covers a wide variety of conformers in the distribution. In addition, it is important to note that MCF does not make any assumptions about the intrinsic geometric factors in conformers like torsional angles and thus provides a simple recipe to scale up the model. With growing number of parameters, larger MCF constantly achieves better performance than smaller counterpart in all metrics. In particular, when compared with MCF-S, MCF-B shows approximately 15% improvement on precision and even larger MCF-L improves it by approximately 20%. The experimental results demonstrate the power of scaling up proposed MCF in better solving conformer generation problem. Since our proposed method simply operates on 3D atomic positions, it provides a straightforward recipe for scaling up the model. This sheds light on how scaling law could potentially benefit applications of deep generative models to scientific domains.
In Fig. 4, we further show a breakdown of the performance on GEOM-DRUGS of MCF with different sizes vs. Torsional diffusion (Jing et al., 2022) as a function of the threshold distance, as well as a function of the number of atoms in molecules. MCF outperforms Torsional Diff. across the full spectrum of thresholds in both recall and precision. When looking at the break-down AMR on different number of atoms in Fig. 4(c), MCF also demonstrates its superior performance for molecules of different sizes. It is indicated that MCF better captures the fine intrinsic geometric structure of conformers and scaling up the model helps improve the performance of proposed model. Also, as the number of parameters increases, MCF demonstrates better performance across all threshold levels in terms of both recall and precision. This provides further evidence on the performance gain from increasing the models size of MCF which is designed to be scalable in a straightforward way. We further investigate the ensemble properties of generated conformers in Appendix A.3.3. Fig. 10 in the Appendix shows examples of MCF generated conformers in GEOM-DRUGS.
5.3 Generalization to GEOM-XL
We now turn to the task of evaluating how well a model trained on GEOM-DRUGS transfers to unseen molecules with large numbers of atoms. Following Jing et al. (2022), we use the GEOM-XL dataset, a subset of GEOM-MoleculeNet that contains 102 molecules with more than 100 atoms. Note that this evaluation not only tests the capacity of models to generalize to larger and more complex molecules but also serves as an out-of-distribution generalization experiment.
@@ 5.3 Generalization to GEOM-XL
MCF-S 2.02 1.87 2.9 2.69 77
MCF-B 1.71 1.61 2.69 2.44 77
MCF-L 1.64 1.51 2.57 2.26 77
Table 3: Generalization results on GEOM-XL.
In Tab. 3 we report AMR for both precision and recall and compare with GeoDiff (Xu et al., 2022), GeoMol (Ganea et al., 2021) and Torsional Diff. (Jing et al., 2022). In particular, when taking the numbers directly from Jing et al. (2022), MCF-B achieves better or comparable performance than Torsional Diff. Further, in running the checkpoint provided by Torsional Diff. and following their validation process we found that 25 molecules failed to be generated, this is due to the fact that Torsional Diff. generates torsional angles conditioned on the molecular graph $\mathcal{G}$ and the local structures obtained from RDKit. And RDKit can fail to find local structures and Torsional Diff. cannot generate conformers in these cases. In our experiments with the same 77 molecules in GEOM-XL from our replica, MCF surpasses Torsional Diff. by a large margin. Leveraging little or no inductive bias in modeling molecular conformers, our proposed method is adaptable to wider variety of data.
Results also show that scaling up the model size to MCF-L further improves the generalizability to large and unseen molecules in GEOM-XL. In our replica, MCF-L demonstrates better performance than smaller model counterparts (i.e.  MCF-S and MCF-B) and surpasses Torsional Diff. by a large margin. The results highlight the generalizability of MCF to large and complex molecules, which may shed a light on pre-training molecular conformer generation model.
5.4 Why does MCF generalize?
A natural question to ask is why does MCF generalize given its non-equivariant design. To answer this question we devised an experiment to understand if conformers in training and validation sets share a canonical coordinate system. In our experiment we apply different rotation transformations to GEOM-QM9 and train MCF on this transformed training set, while keeping the validation set unchanged. Three rotation transformations are investigated: 1) “Fixed” applies a single random rotation to all conformers, 2) “Variable” applies a different rotation to each conformer and keeps it through training, 3) “Random” applies a different random rotation to each conformer in each training epoch. Fig. 5 shows the results in these different settings.
Refer to caption Refer to caption
(a) (b)
Figure 5: (a) Mean Coverage and (b) mean AMR of different rotation augmentation strategies on GEOM-QM9 when compared with training on original dataset.
Not surprisingly, applying a fixed rotation to the training set minimally affects performance. This is because a fixed rotation does not break relative SO(3) relations between conformers in the training set. However, rotating each conformer independently once during training (e.g. “Variable”) negatively impacts performance. This finding points to the fact that the DFT simulations used to generate the data might be implicitly encoding a canonical coordinate system, which affects generalization if broken. Finally, applying a random rotation to each conformer on each training epoch forces the model to be invariant to any coordinate system, which is a more challenging task. Notably, though randomly rotating each conformer leads to worse results in recall, the performance drop is still marginal. Finally, by training a bigger model on this randomly rotated training set (i.e.  Random-L) we can recover most of the performance gap in comparison with training on the original dataset.
These experiments show that inductive biases like roto-equivariance can be traded for scale in general-purpose models. Our results highlight that a domain-agnostic model at scale can achieve better performance than intricate models with strong domain-specific inductive biases in molecular conformer generation. We hope our findings will inspire the community to develop simple models that are prone to benefit from “scaling laws” especially when taking into account the fast growth of available scientific data.
5.5 Sampling
Refer to caption Refer to caption
(a) (b)
Refer to caption Refer to caption
(c) (d)
Figure 6: Inference wall clock time v.s. (a) precision coverage, (b) recall coverage, (c) precision AMR, and (d) recall AMR with Torsional Diff. and our MCF.
In this section, we investigate the performance of our MCF under limited computation budget in inference. To this end, we report COV and AMR of MCF with respect to different wall-clock sampling times. DDIM (Song et al., 2021), an efficient sampler, is applied, which uses a significantly smaller number of sampling steps than vanilla DDPM (i.e.  1000 sample steps). Specifically, we sample conformers with 3, 5, 10, 20, and 50 sampling steps with DDIM and compare the performance as well as inference time with Torsional Diff. (Jing et al., 2022). All models are benchmarked on a single A100 GPU for comparison. It is shown that MCF is more efficient than Torsional Diffusion, which means that for the same inference time in seconds, MCF always outperforms Torsional Diffusion across all metrics and all model sizes. Notably, due to application of equivariant operations, Torsional Diff. can be time consuming in inference. Even when using a very limited sampling steps (e.g. 5 steps) MCF achieves comparable or better COV and AMR than Torsional Diff. using 50 sampling steps, while being more efficient in wall-clock inference time. MCF with different sizes achieves Pareto frontier of performance (i.e.  COV and AMR) and sampling efficiency compared with Torsional Diff. This further indicates rotational or translational equivariance may not be a strong requirement while simple and scalable framework like MCF can own the merits in efficiency. Examples of MCF sampled conformers with different sampling steps can be found in Fig. 9 in the Appendix.